Please help:
Find the equation of a line with parallel to y=8x-13 with y-intercept 9


Thanks,

Answer by fractalier(2101)

Parallel means having the same slope...
The slope of y = 8x - 13 is 8.
Since it goes thru (0, 9), we can use the point-slope form:
y - y1 = m(x - x1)
y - 9 = 8(x - 0)
y = 8x + 9